OpenStudy (anonymous):

Find the standard form: Focus: (4, -9) Directrix: x=2

OpenStudy (anonymous):

I'm guessing the question is asking the standard form of a "parabola" when given this information. You're given that the focus has coordinates of: \[\large S(4, -9)\] and the equation of the Directrix: \[\large D:x=2\] Okay, the vertex of the parabola is in between the directrix and the focus. So the coordinates of the vertex would be (3, -9) Okay. The general formula for this "side-ways parabola" would be: \[\large (y-k)^2=4a(x-h)\] where:(h, k) is the vertex. And a is the length between the vertex and the directrix/focus. It could of been -4a(x-h) but as you can see, the focus is to the right of the directrix which means that it must be a "positive side-ways parabola". The value of a would be 1 since the vertex is 1 unit apart from the directrix and the focus. So now we have all the values we need. The standard form of that parabola would be: \[\large (x-(3))^2=4(1)[y-(-9)]\] \[\large (x-3)^2=4(y+9)\]